# Risk-Aware Autonomous Driving with Linear Temporal Logic Specifications

### v2

@@ Abstract
Humans naturally balance the risks of different concerns while driving, including traffic rule violations, minor accidents, and fatalities. However, achieving the same behavior in autonomous systems remains an open problem. This paper extends a risk metric that has been verified in human-like driving studies to encompass more complex driving scenarios specified by linear temporal logic (LTL) that go beyond just collision risks. This extension incorporates the timing and severity of events into LTL specifications, thereby reflecting a human-like risk awareness. Without sacrificing expressivity for traffic rules, we adopt LTL specifications composed of safety and co-safety formulas, allowing the control synthesis problem to be reformulated as a reachability problem. By leveraging occupation measures, we formulate a linear programming (LP) problem for this LTL-based risk metric. Consequently, the synthesized policy balances different types of risks, including not only collision risks but also traffic rule violations. The effectiveness of the proposed approach is validated by three typical traffic scenarios in the Carla simulator.
@@ I Introduction
Developing motion planning frameworks for autonomous vehicles remains a significant challenge due to the inherent complexity and uncertainty of real-world traffic [1, 2]. The complexity arises primarily from the diverse concerns in practical driving tasks, including safety [3], traffic rules [4], and social norms [5]. The uncertainty mainly originates from the dynamic environment with unpredictable traffic participants [6] and unexpected events [7]. Taking an unprotected turn scenario in Fig. 1 as an example, the ego vehicle must make a left turn while avoiding collisions with the opponent car, obey traffic rules such as reacting to traffic lights, and follow driving norms like yielding to oncoming vehicles. Environmental uncertainty imposes the ego vehicle at the potential risk of violating the above specifications. Therefore, a risk-aware autonomous driving scheme is essential to make decisions that balance diverse risks associated with safety, traffic rules, and social norms [8].
@@
Temporal logic provides a formal and expressive framework for describing high-level tasks that incorporate spatial and temporal constraints and synthesize provably correct controllers [9, 10, 11], which makes it well suited for encoding traffic tasks with rules and norms [12] and improving trustworthiness in motion planning tasks [13, 14]. While temporal logic-based planning offers significant advantages, existing frameworks remain limited and impractical for real-world traffic deployment. Due to uncertainty and dynamic interaction in real traffic, achieving complete safety and full satisfaction of temporal logic specifications is unrealistic, making it essential to develop a risk-aware framework. Existing temporal logic planning methods quantify risk as the probability that a temporal logic specification fails and focus on maximizing the satisfaction probability of these specifications [15, 16, 17, 18, 19]. However, such risk metrics are not suitable for realistic traffic where different types of risky events may be involved [20]. In practical traffic with diverse risks, human drivers intuitively prioritize near-future and more severe events over those in the distant future or with less severity [21]. Unfortunately, the satisfaction probability of temporal logic specifications fails to differentiate the severity and timing of events [22], which results in illogical and even unsafe driving behavior.
@@
The driver’s risk field model [23, 24], proposed by the transportation community, has been shown to align with human-like risk awareness by discounting collision risks over the prediction horizon. This mode enables vehicles to track references safely by balancing risks between multiple obstacles. However, this model is limited to collision avoidance in tracking tasks, making it less effective for handling realistic traffic scenarios with complex rules. To address this, we extend the risk field model by integrating temporal logic specifications. This paper selects to use LTL [9], a classic type of temporal logic widely used in motion planning. Specifically, we extend upon the tracking task with more complex driving goals expressed by co-safe specifications and quantify risk specified as breaching safety specifications. By incorporating the inherent ability to balance timing and severity from the risk field model, this extension overcomes the limitation of temporal logic mentioned above. Safety or co-safety specifications [25] are special fragments of LTL that can transform the control synthesis problem into a classical reachability problem. Therefore, this approach retains the ability to simplify the LTL planning into a reachability problem while preserving the necessary expressivity for traffic tasks. To formulate a tractable control synthesis problem, we rewrite the risk metric using the discounted version of occupation measures, a key concept for solving a stochastic reachability problem for a Markov Decision Process (MDP) [26]. Referring to the multi-objective control synthesis formulation for LTL in [27], we formulate a linear programming (LP) problem over product automaton to synthesize risk-aware policies using occupation measures as decision variables.
@@
The effectiveness and scalability of the proposed method are validated in three distinct driving scenarios using the high-fidelity Carla simulator [28], demonstrating its capability to handle complexities from traffic rules and unexpected events due to uncertainty, and its scalability across versatile scenarios.
@@ II-B Specifications: Linear Temporal Logic (LTL)
[…]
@@
Automaton: A general LTL specification can be translated into a Büchi automaton [9] whose semantics are defined over infinite words. In contrast, scLTL and safety specification can be also translated into deterministic finite-state automatons (DFA), defined as follows.
@@ II-C Risk-Aware Control Synthesis Problem
Since traffic rules are primarily concerned with the occurrence of good or bad events rather than repetitive behaviors, we use the combination of safety and co-safety fragments to effectively represent nearly all traffic tasks and rules while mitigating the computational demand of control synthesis. Thus, we consider an LTL specification in the form of $\psi=\psi_{cs}\wedge\psi_{s}$ defined over an atomic proposition set $\mathsf{AP}$, composed with a safety formula $\psi_{s}$ and a co-safety formula $\psi_{cs}$. The corresponding DFAs of $\psi_{cs}$ and $\psi_{s}$ are denoted by $\mathcal{A}_{cs}$ and $\mathcal{A}_{s}$, respectively. To represent the uncertain behavior of both the environment and vehicle, we use a composed MDP $\overline{\mathcal{M}}$, as defined in Definition 1. This composed MDP combines an MDP model for an ego vehicle, $\mathcal{M}_{v}$ defined in Def. 1, and an MC model for an uncontrollable environment, $\mathcal{C}_{e}$ defined in Definition 3. The states of $\overline{\mathcal{M}}$ are labeled by $\mathcal{L}$ to an alphabet set $\Sigma$, which can be used to verify the satisfaction of $\psi$. To handle the uncertainty in realistic traffic like a human driver, we aim to develop a human-like risk metric for $\psi$ and then synthesize a stochastic policy $\pi\!\in\!\Pi$ that respects a human-like risk metric.
@@
Risk metrics are often defined based on simple reach-avoid problems [32]. Therefore, we will start by rewriting the satisfaction problem for the given specifications $\psi=\psi_{cs}\wedge\psi_{s}$ as a reach-avoid problem over product MDP.
@@ III-A Product MDP and Reachability Problem
It is well known [9] that the satisfaction probability of co-safety specification equates to a reachability probability towards the accepting state set of the product automaton defined in Def. 5,
@@
where $\mathbf{1}(\cdot)$ is an indicator function, $\bar{\pi}$ is policy for $\mathcal{P}$, and ¯$\overline{\mathbb{P}}_{\pi}$ denotes the probability distribution over the paths of $\mathcal{P}$ under $\bar{\pi}$. Not that $\bar{\pi}$ for $\mathcal{P}$ that is a Markovian and stationary policy, which is converted from $\pi$ for $\mathcal{M}_{v}$ defined in Def. 2. Similarly, for every safety specification $\psi_{s}$, the violation probability equates to a reachability probability towards the non-accepting state set of $\mathcal{P}$,
@@ III-B Limitations of Event Probability
Using satisfaction and violation probability defined in (2) and (3) to evaluate safety in realistic traffic might lead to unnatural behavior. The event probability in (2) and (3) means that each possible trace generated by $\pi$ that drives $\mathcal{P}$ to enter into final states in finite time gets weight 1 and all else gets weighted 0, which is displayed in Fig. 2. Such an approach embraces two points against the awareness way of human drivers: (1) It does not account for the timing of violations, which may cause the vehicle to react too early to dangers that happen in the distant future; (2) It fails to consider varying costs associated with various types of violations, preventing the vehicle from appropriately balancing the seriousness of various dangers. Thus, a risk metric reflecting human-like risk awareness is necessary.
@@ III-C Human-like Metric Design
Discounted rewards are commonly used in reinforcement learning and economics to account for the timing of events when decisions lead to rewards at different moments in the future. The satisfaction probability, $\mathbb{P}_{\pi}(\boldsymbol{\omega}\models\psi_{cs})$, can be modified by introducing discounting as follows,
@@
where $\gamma<1$ is a discount factor. This discounting represents the preference of people for obtaining rewards sooner rather than later. The rationality of the discounting operation has also been supported by psychology studies [33]. When it comes to violation risks of safety specification, we propose a similar metric. In addition, varying severity of events are considered by a cost mapping function, denoted as $c(\cdot)\!:\!Z\!\rightarrow\!\mathbb{R}$, that assigns cost values to different product states. Specifically, $c(\hat{z})=0$ if $\hat{z}\!\notin\!D$, while $c(\hat{z})>0$ if $\hat{z}\!\in\!D$. Furthermore, we propose the following metric to evaluate the risk of a policy $\pi$,
@@
Figure 3: Sketch for $\mathcal{R}_{\bar{\pi}}$. The grey arrows indicate the awareness levels assigned to violation events, while the colored arrows represent the corresponding risk values. Red and blue colors differentiate between two different types of events.
{mdframed}
Driver’s Risk Field (DRF) model is designed empirically to evaluate the uncertainty distribution from a driver’s view. The risk field $\alpha(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ maps horizontal coordinates to scalar values representing the driver’s attention. To model human-like awareness, DRF is designed based on the intuition that the importance of future events diminishes while the corresponding uncertainty increases along the predictive horizon. A detailed formulation of DRF can be found in [23]. Perceived risk is evaluated by calculating the expected event cost, which is the product of the risk field and the cost map. With a grid discretization, the DRF risk metric is defined as:
$\mathcal{R}=\sum_{x\in X}\sum_{y\in Y}c(x,y)\alpha(x,y),$ (6)
where $c(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ represents the cost distribution over horizontal coordinates, and $X$ and $Y$ are finite sets of discretized coordinates along with $x$ and $y$-axis.
@@
Referring to the standard occupation measure formulated as $\mathbb{E}_{\bar{\pi}}\textstyle\left[\sum_{t}^{\infty}\mathbf{1}(z_{t}=z)\right]$, $\beta_{\bar{\pi}}(z)$ can be regarded as a discounted version of the occupation measure. An occupation measure describes the expected number of times that a state has been visited before reaching the final set, given a Markove and stationary policy $\bar{\pi}(z,a)$. Occupation measures can be used to formulate stochastic shortest path problems [34]. By comparing risk metrics in (7) and (6), it is evident that the discounted occupation measure plays a similar role as the DRF model. In addition, the discounted occupation measure embraces similar awareness for events as the DRF model. Thus, we conjecture that the proposed risk metric embodies a similar human-like risk awareness as DRF.
@@ IV Risk-aware control
In this section, we formulate a first risk-aware control synthesis problem with the proposed metrics. A reasonable approach is to maximize the rewards metric of co-safety specification and also keep the risk metric of safety specification in an acceptable range, which is formulated as follows,
@@
where $r_{th}\in\mathbb{R}$ is a constant threshold. We first reformulate this optimization problem into a linear programming problem. Afterward, we detail how traffic rules and vehicle dynamics should be integrated and evaluate the risk metric in a traffic scenario.
@@ IV-A Linear Programming Problem
The optimal policy $\bar{\pi}$ can be extracted from the optimized occupation measures using Eq. (11), and then $\bar{\pi}$ can convert to $\pi\!\in\!\Pi$ defined on $\mathcal{M}_{v}$. This LP problem can be solved using an off-the-shelf solver, Gurobi [35].
To visualize the risk field calculated by the LP problem intuitively, we first consider a reach-avoid problem over the 2-dimensional positions of the ego vehicle, excluding the product states that include environment states and specification automaton states. Fig. 4 displays the risk field with discounted occupation measures obtained by solving the LP problem in Eq. (13). The red solid line represents the optimal path determined by the optimized risk field. As shown in the figure, we can see that the risk field obtained by the occupation measure aligns with the pattern of the proposed risk field in [23], revealing the underlying equivalence between the proposed metric and DRF. In addition, the different risk thresholds $r_{th}$ in Eq. (13) will alter the shape of the risk field and further tune the conservatism of the generated policy.

### v3

@@ Abstract
Human drivers naturally balance the risks of different concerns while driving, including traffic rule violations, minor accidents, and fatalities. However, achieving the same behavior in autonomous driving systems remains an open problem. This paper extends a risk metric that has been verified in human-like driving studies to encompass more complex driving scenarios specified by linear temporal logic (LTL) that go beyond just collision risks. This extension incorporates the timing and severity of events into LTL specifications, thereby reflecting a human-like risk awareness. Without sacrificing expressivity for traffic rules, we adopt LTL specifications composed of safety and co-safety formulas, allowing the control synthesis problem to be reformulated as a reachability problem. By leveraging occupation measures, we further formulate a linear programming (LP) problem for this LTL-based risk metric. Consequently, the synthesized policy balances different types of driving risks, including both collision risks and traffic rule violations. The effectiveness of the proposed approach is validated by three typical traffic scenarios in Carla simulator.
@@ I Introduction
Developing motion planning frameworks for autonomous vehicles remains a significant challenge due to the inherent complexity and uncertainty of real-world traffic [1]. The complexity arises primarily from the diverse concerns in practical driving tasks, including safety, traffic rules, and social norms [2]. The uncertainty mainly originates from the dynamic environment with unpredictable traffic participants and unexpected events [3]. Taking an unprotected turn scenario in Fig. 1 as an example, the ego vehicle must make a left turn while avoiding collisions with the opponent car, obey traffic rules such as reacting to traffic lights, and follow driving norms like yielding to oncoming vehicles. Environmental uncertainty places the ego vehicle at the potential risk of violating the above specifications. Therefore, a risk-aware autonomous driving scheme is essential to make decisions that balance diverse risks associated with safety, traffic rules, and social norms.
@@
Temporal logic is a formal and expressive language used to describe both spatial and temporal properties of systems through logical and temporal operators [4]. It has been widely applied in verifying software and digital circuits, and more recently explored in motion planning of autonomous vehicles, as it is well-suited for encoding high-level traffic rules [5] and synthesizing provably correct controllers [6, 7]. However, current temporal logic-based planning frameworks remain limited and impractical for traffic applications due to inherent uncertainty in the real world. Conventional temporal logic planning aims to achieve complete safety and full satisfaction of temporal logic specifications [7], which is unrealistic in practice. To handle uncertainty, existing methods quantify risk using the probability that a temporal logic specification fails and focus on maximizing the satisfaction probability of these specifications [8, 9, 10]. However, such risk metrics are not suitable for realistic traffic where different types of risky events are involved [11]. In real-world traffic with diverse risks, human drivers intuitively prioritize near-future and more severe events over those in the distant future or with less severity [12]. Unfortunately, the satisfaction probability of temporal logic fails to differentiate the severity and timing of multiple events, which may result in illogical and even unsafe behavior.
@@
To achieve human-like risk assessment, the driver’s risk field model is proposed in the transportation domain, leveraging an expert-designed distribution to emulate human drivers’ awareness of predictive events [13, 14]. This heuristic field model enables vehicles to differentiate the timing and severity of potential collisions and track reference paths safely by balancing risks among multiple obstacles. However, this model is limited to collision avoidance in tracking tasks, making it less effective for handling realistic traffic scenarios with complex rules. To address this, we extend the risk field model to temporal logic specifications. This paper uses linear temporal logic (LTL) [4], a classic type of temporal logic widely used in motion planning. By incorporating the inherent ability to balance timing and severity from the risk field model, this extension overcomes the limitation of temporal logic mentioned above. Specifically, we extend upon the tracking task with more complex driving goals expressed by co-safe specifications and quantify risk specified as violating safety specifications. Safety or co-safety specifications [15] are special fragments of LTL that can transform the control synthesis problem into a classical reachability problem. Therefore, this approach reduces the LTL planning into a reachability problem while preserving the necessary expressivity for traffic tasks. To formulate a tractable control synthesis problem, we rewrite the risk metric using the discounted version of occupation measures, a key concept used in the stochastic reachability problem for constrained Markov Decision Process (MDP) models [16]. Referring to the multi-objective control synthesis formulation for LTL in [17], we formulate a linear programming (LP) problem over a product automaton to synthesize risk-aware policies using occupation measures as decision variables.
@@
The main contributions of this paper are the introduction of a human-like LTL-based risk metric and the development of a corresponding risk-aware controller to balance different types of risks. The effectiveness and scalability of the proposed method are validated in three distinct driving scenarios on Carla simulator [18], demonstrating its capability to handle complexities from traffic rules and unexpected events due to uncertainty, and its scalability across versatile scenarios.
@@ II-B Specifications: Linear Temporal Logic (LTL)
Semantics: For a given word $\boldsymbol{\omega}$, basic semantics of LTL are given as $\boldsymbol{\omega}_{k}\!\models\!p$, if $p\in\omega_{k}$; $\boldsymbol{\omega}_{k}\!\models\!\lnot p$, if $p\notin\omega_{k}$; $\boldsymbol{\omega}_{k}\!\models\!\psi_{1}\wedge\psi_{2}$, if $\boldsymbol{\omega}_{k}\!\models\!\psi_{1}$ and $\boldsymbol{\omega}_{k}\!\models\!\psi_{2}$; $\boldsymbol{\omega}_{k}\!\models\!\bigcirc\psi$, if $\boldsymbol{\omega}_{k+1}\!\models\!\psi$; $\boldsymbol{\omega}_{k}\!\models\!\psi_{1}\mathsf{U}\psi_{2}$, if $\exists$ $i\in\mathbb{N}$ such that $\boldsymbol{\omega}_{k+i}\!\models\!\psi_{2}$, and $\boldsymbol{\omega}_{k+j}\!\models\!\psi_{1}$ holds  $\forall\,0\!\leq j\!<\!i$.
@@
Automaton: A general LTL specification can be translated into a Büchi automaton [4] whose semantics are defined over infinite words. In contrast, scLTL and safety specification can be also translated into deterministic finite-state automatons (DFA), defined as follows:
@@ II-C Risk-Aware Control Synthesis Problem
Since traffic rules are primarily concerned with the occurrence of good or bad events rather than repetitive behaviors, we use the combination of safety and co-safety fragments to effectively represent nearly all traffic tasks and rules while mitigating the computational demand of control synthesis. Thus, we consider an LTL specification in the form of $\psi=\psi_{cs}\wedge\psi_{s}$ defined over an atomic proposition set $\mathsf{AP}$, composed with a safety formula $\psi_{s}$ and a co-safety formula $\psi_{cs}$. The corresponding DFAs of $\psi_{cs}$ and $\psi_{s}$ are denoted by $\mathcal{A}_{cs}$ and $\mathcal{A}_{s}$, respectively. To represent the uncertain behavior of both the environment and vehicle, we use a composed MDP $\overline{\mathcal{M}}$, as defined in Def. 1. This composed MDP combines an MDP model for an ego vehicle, $\mathcal{M}_{v}$, and an MC model for an uncontrollable environment, $\mathcal{C}_{e}$, defined in Def. 3. The states of $\overline{\mathcal{M}}$ are labeled by $\mathcal{L}$ to an alphabet set $\Sigma$, which can be used to verify the satisfaction of $\psi$. To handle the uncertainty in realistic traffic like a human driver, we aim to develop a human-like risk metric for $\psi$ and then synthesize a policy $\pi\!\in\!\Pi$ that balances different types of risk and drives the ego vehicle to satisfy $\psi$.
@@
Risk assessments are often discussed on simple reach-avoid problems [22]. Therefore, we start by rewriting the satisfaction problem for the given specifications $\psi=\psi_{cs}\wedge\psi_{s}$ as a reach-avoid problem over product MDP.
@@ III-A Product MDP and Reachability Problem
It is well known from [4] that the satisfaction probability of a co-safety specification equates to a reachability probability towards the accepting state set of the product automaton defined in Def. 5,
@@
where $\mathbf{1}(\cdot)$ is an indicator function, $\bar{\pi}$ is policy for $\mathcal{P}$, and ¯$\overline{\mathbb{P}}_{\pi}$ denotes the probability distribution over the paths of $\mathcal{P}$ under $\bar{\pi}$. Note that $\bar{\pi}$ for $\mathcal{P}$ is a Markovian and stationary policy, which is converted from $\pi$ for $\mathcal{M}_{v}$ defined in Def. 2. Similarly, for a safety specification $\psi_{s}$, the violation probability equates to a reachability probability towards the non-accepting state set of $\mathcal{P}$,
@@ III-B Limitations of Event Probability
The event probability in (2) and (3) means that each possible trace generated by $\pi$ gets weight 1 if it enters into final states, and otherwise gets weighted 0. As displayed in Fig. 2, three different traces are considered identical, even when they enter the final set at different times. Such an approach embraces two points against the awareness way of human drivers: (1) It does not account for the timing of violations, which may cause the vehicle to react too early to dangers that happen in the distant future; (2) It fails to distinguish different types of events, preventing the vehicle from appropriately account for the severity of multiple risks. Thus, a new risk metric for LTL specifications reflecting human-like awareness is necessary.
@@ III-C Human-like Metric Design
Discounted rewards are commonly used in reinforcement learning and economics to account for the timing of events when decisions lead to rewards at different future moments. The satisfaction probability, $\mathbb{P}_{\pi}(\boldsymbol{\omega}\models\psi_{cs})$, can be modified by introducing discounting as follows,
@@
where $\gamma\!<\!1$ is a discount factor. This discounting represents the preference of people for obtaining rewards sooner rather than later. The rationality of the discounting operation has also been supported by psychology studies [23]. When it comes to violation risks of safety specifications, we propose a similar metric. In addition, varying severity of events is considered by a cost mapping function, denoted as $c(\cdot)\!:\!Z\!\rightarrow\!\mathbb{R}$, that assigns cost values to different product states. Specifically, $c(\hat{z})=0$ if $\hat{z}\!\notin\!D$, while $c(\hat{z})>0$ if $\hat{z}\!\in\!D$. Furthermore, we propose the following metric to evaluate the risk of a policy $\pi$,
@@
Figure 3: Sketch for $\mathcal{R}_{\bar{\pi}}$. Red and blue indicate two types of events with different severity levels. The grey arrows illustrate decreasing awareness levels over the horizon due to discounting, while the colored arrows represent the corresponding risk values.
Driver’s Risk Field (DRF) is a heuristic model to evaluate the risks from a driver’s view. The risk field $\alpha(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ maps planar coordinates to scalar values representing the driver’s attention. To model human-like awareness, DRF is designed based on the intuition that the importance of future events decreases while the corresponding uncertainty increases along the predictive horizon. A detailed formulation of DRF can be found in [13]. Perceived risk is evaluated by calculating the expected event cost, which is the product of the risk field and the cost map. With a grid discretization, the DRF risk metric is defined as: $\mathcal{R}=\sum_{x\in X}\sum_{y\in Y}c(x,y)\alpha(x,y),$ (6) where $c(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ represents the cost distribution over planar coordinates, and $X$ and $Y$ are finite sets of discretized coordinates along with $x$ and $y$-axis.
@@
Referring to the standard occupation measure defined as $\mathbb{E}_{\bar{\pi}}\textstyle\left[\sum_{t}^{\infty}\mathbf{1}(z_{t}=z)\right]$, $\beta_{\bar{\pi}}(z)$ can be regarded as a discounted version of the occupation measure. An occupation measure describes the expected number of times that a state has been visited before reaching the final set, given a Markov and stationary policy $\bar{\pi}(z,a)$. Occupation measures can be used to formulate stochastic shortest path problems [24]. By comparing risk metrics in (7) and (6), it is evident that the discounted occupation measure plays a similar role as the DRF model. In addition, the discounted occupation measure embraces similar awareness for events as the DRF model. Thus, we conjecture that the proposed risk metric captures a similar human-like risk awareness as DRF, which will be validated by subsequent simulation results.
Remark 1.
The cost function is designed by experts in this work but can potentially be learned from real-world driving data to emulate human awareness of event severity.
@@ IV Risk-aware control
In this section, we first formulate the control synthesis problem with the proposed risk metric. A natural approach is to maximize the rewards metric of co-safety specification and also keep the risk metric of safety specification in an acceptable range, as formulated below:
@@
where $r_{th}\in\mathbb{R}$ is a constant threshold. We first reformulate this optimization problem into an LP problem. Afterward, we detail the control synthesis framework that integrates traffic rules and vehicle dynamics, followed by an evaluation of the synthesized policy in a simplified traffic scenario.
@@ IV-A Linear Programming Problem
The optimal policy $\bar{\pi}$ can be extracted from the optimized occupation measures using Eq. (11), and then $\bar{\pi}$ can convert to $\pi\!\in\!\Pi$ defined on $\mathcal{M}_{v}$. This LP problem can be solved using Gurobi [25] solver.
To visualize the risk field calculated by the LP problem intuitively, we first consider a reach-avoid problem over the planar position space of a vehicle, excluding the dynamic environment and LTL specifications. Fig. 4 displays the risk field comprising discounted occupation measures obtained by solving the LP problem in Eq. (13). The red solid line represents the optimal path determined by the occupation measures. As shown in the figure, we can see that the risk field obtained by the occupation measure aligns with the pattern of the heuristically designed risk field in [13], revealing the underlying equivalence between the proposed metric and DRF. In addition, the different risk thresholds $r_{th}$ in Eq. (13) will alter the shape of the risk field and further tune the conservatism of the generated policy.